# Weather4Cast 2025

### v1

@@ 1 Introduction
Extreme precipitation poses outsized risks to life, property, and commerce. Recent studies predict that these events are likely to increase in frequency and severity during the coming decades, spurred on by a warming atmosphere [Madakumbura_2021, Kotz_2022]. Timely and accurate public advisories save lives and allow interested parties to make appropriate preparations for future weather events. However, traditional sources of operational forecast guidance are plagued by key limitations. First, numerical weather prediction (NWP) models like the operational ECMWF incur enormous computational costs. And second, these models are often only available at hourly time-scales, and are therefore inappropriate sources of real-time guidance.
Recent data-driven, so-called AI-weather prediction (AI-WP) approaches offer a compute-efficient alternative to NWP models by learning to emulate forecast analysis data. Breakthrough works from Google [Lam2023, Price2024], NVIDIA [pathak2022fourcastnetglobaldatadrivenhighresolution], Huawei [Bi_2023], and others replace core components of the weather-prediction pipeline with neural networks. These learning-based methods have manifold advantages. Namely, AI-WP pipelines can practically run at sub-hourly intervals and in large ensembles; researchers are therefore optimistic that operational meteorologists can one day leverage these models as a source of real-time guidance and robust uncertainty estimation. However, previous studies primarily explore AI-WP for medium-range forecasting on a global scale. This is a natural choice given the surplus of historical data available at these resolutions. By contrast, forecasting extreme precipitation at nowcasting timescales is a nascent area of research.
@@ 2.1 Weather4Cast competition
Beginning in 2021, the annual Weather4Cast competition (W4CC) evaluates machine learning (ML) techniques for weather forecasting within spatial and temporal transfer learning regimes [pmlr-v220-gruca23a, gruca2024multitask]. Each year, participants are tasked with developing approaches to predict unseen (OPERA) high-resolution radar precipitation fields (or targets derived from these data) using 11-band HRIT geostationary satellite data and static topographic maps as context. Competition data cover 10, non-overlapping regions (spatial areas of $\times1512$ km) across Europe, northern Africa, and the middle-east for which HRIT data are completely available, while OPERA data are sparse.
@@ 3.1 W4CC 2021
In its inaugural year organizers proposed a multivariate weather forecasting challenge: participants predicted future atmospheric states (represented by four variables) for eight hours of lead-time using the same features and topographic information as model inputs [kwok2021variationalunetweatherforecasting]. The authors of [kwok2021variationalunetweatherforecasting] proposed to solve this task by training a U-Net VAE with a combination of deterministic and probabilistic losses. In this work it is argued that these design choices are crucial for modeling the fundamental uncertainty (i.e., aleatoric and epistemic) present in weather forecasting and that they therefore aid model performance. However, this complex, multi-loss setup may prove difficult to extend to new tasks in practice.
@@ 3.2 W4CC 2022
In [belousov2022solvingweather4castchallengevisual] an ensemble of video classification models (e.g., Swin UNETR, ViViT) is presented [hatamizadeh2022swinunetrswintransformers, arnab2021vivitvideovisiontransformer]). An obvious drawback is that this approach requires training three different models from scratch for a single task and incurs further computational costs by bootstapping predictions at inference-time. Additionally, modifications made by the authors to adopt video classification models to a video generation task (e.g., appending a single convolutional layer to ViViT to predict the output temporal dimension) are likely brittle design choices. In [li2022superresolutionprobabilisticrainprediction], the authors apply a 3D U-Net and an Earthformer [çiçek20163dunetlearningdense, gao2023earthformerexploringspacetimetransformers] to solve this year’s stage 1 and stage 2 challenges, respectively. In the second challenge, an Earthformer is employed to efficiently decompose a video-sequence into cuboids (as opposed to frame-wise patches as proposed in ViT); self-attention is performed between cuboids and a collection of global context tokens $\mathcal{G}$. However, inter-cuboid attention is axial, and therefore, cuboids only learn spatial dependencies from a set of coarse global context vectors.
Beginning with a standard 3D U-Net, the authors of [park2022rainunetsuperresolutionrainmovie] introduce a time-space (TS) convolutional block in which a spatial convolution, followed by a “temporal-wise dilated convolution” (TSDConv), and finally a temporal convolution are applied to the input feature maps. The authors note that their proposed TSDConv is simply an extension of the large kernel convolution operation into three dimensions. Additionally, WeatherFusionNet [pihrt2022weatherfusionnetpredictingprecipitationsatellite] propose a physically informed U-Net-based architecture. Their model as inputs uses HRIT satellite radiances, a PDE-constrained video prediction via PhyDNet [guen2020disentanglingphysicaldynamicsunknown], and the outputs of a proposed Sat2Rad module to forecast binary rain masks. While the methods above achieve strong performance on W4CC tasks, most are fundamentally reliant on CNN model designs, and likely suffer from the inductive biases and poor scaling these architectures exhibit in other domains.
@@ 3.3 W4CC 2023
This year, researchers at Alibaba Cloud proposed a mixture-of-experts-inspired-solution to both the “core” and “transfer learning” competition tracks [li2023precipitationpredictionusingensemble]. The authors design a “predictor”, which samples multiple times from a WeatherFusionNet backbone, and a “controller”; the latter module produces a final output as a weighted combination of the expert predictions. In addition to being a minor extension of prior works, the proposed model employs an arduous three-stage training scheme. The aforementioned work is still preferable to [kumar2023skilfulprecipitationnowcastingusin], who simply finetune a NowcastNet [Zhang2023] on the challenge dataset. In [han2023learningrobustprecipitationforecaster] an augmentation technique is proposed to sample linear interpolations of the original training samples. By formulating precipitation forecasting as a multi-class, video classification task (where each per-pixel class represents a discrete interval of precipitation, the authors may employ a modified dice-loss with an added ordinal prior (i.e., under-predictions are punished more heavily than over-predictions). While certainly performant, this approach is heavily engineered and therefore may generalize poorly to new tasks and datasets
@@ 3.4 W4CC 2024
[deshpande2024conditionalgenerativeadversarialnetwork] develop a conditional GAN network trained on a sparsely-sampled subset of the original HRIT channels. Here, the authors completely discard near-infrared, water vapor, and visible satallite data, and average the remaining 4 channels; after data preprocessing their model is trained on only single-channel inputs. Further preprocessing techniques are employed to improve model performance: including thresholding out non-cloudy regions and extracting optical flow scores from the input sequence. Similar to [han2023learningrobustprecipitationforecaster], this work proposes a bespoke, highly-engineered pipeline that is unlikely to extend gracefully to future work.
@@ 4.1 SaTformer model
Following prior work [dosovitskiy2021imageworth16x16words, bertasius2021spacetimeattentionneedvideo], we partition each frame in the input sequence into $N=HW/P^{2}$ non-overlapping patches with side-length $P$ indexed at timestep $t\in 1,...,T$ and spatial index $p\in 1,...,N$: $\mathbf{x}^{(0)}_{(t,p)}\in\mathbb{R}^{P\times P\times C}$; patches completely cover the spatial dimensions of the original frame. We flatten each patch, then project the resulting vectors into tokens: $\mathbf{z^{\prime}}^{(0)}_{(t,p)}\in\mathbb{R}^{d}$. We prepend a randomly initialized “class token” $\mathbf{z^{\prime}}_{(0,0)}^{(0)}$ to the token sequence intended to aggregate relevant information for prediction of the target label:
$\mathbf{z^{\prime}}_{(0,0)}^{(0)},\mathbf{z^{\prime}}^{(0)}_{(1,1)},...,\mathbf{z^{\prime}}^{(0)}_{(T,N)}]$ (1)
@@ 4.1.1 Full space-time self-attention
Where $\mathbf{q}^{(l)}_{(t,p)}\in\mathbb{R}^{d}$, $\mathbf{k}^{(l)}_{(t,p)}\in\mathbb{R}^{d}$ $\mathbf{v}^{(l)}_{(t,p)}\in\mathbb{R}^{d}$ and $\text{LN}(\cdot)$ denotes the LayerNorm operation [ba2016layernormalization]. Next we compute importance weights $\mathbf{a}^{(l)}_{(t,p)}$ between all pairs of tokens across time and space:
$\mathbf{a}^{(l)}_{(t,p)}=\text{softmax}\big(\frac{(\mathbf{q}^{(l)}_{(t,p)})^{\top}}{\sqrt{d}}\cdot[\mathbf{k}_{(0,0)}^{(l)},\mathbf{k}^{(l)}_{(1,1)},...,\mathbf{k}^{(l)}_{(T,N)}]\big)$ (4)
@@
$\mathbf{s}^{(l)}_{(t,p)}=\mathbf{a}^{(l)}_{(t,p)(0,0)}\mathbf{v}^{(l)}_{(0,0)}+\sum_{t^{\prime}=1}^{T}\sum_{p^{\prime}=1}^{N}\underbrace{\mathbf{a}^{(l)}_{(t,p)(t^{\prime},p^{\prime})}}_{\text{scaler}}\mathbf{v}^{(l)}_{(t^{\prime},p^{\prime})}$ (5)
@@
$\mathbf{z^{\prime}}^{(l)}_{(t,p)}={s}^{(l)}_{(t,p)}W_{\text{out}}+\mathbf{z}^{(l-1)}_{(t,p)}$ $\mathbf{z}^{(l)}_{(t,p)}=\text{MLP}(\text{LN}(\mathbf{z^{\prime}}^{(l)}_{(t,p)}))+\mathbf{z^{\prime}}^{(l)}_{(t,p)}$ (6)
@@ 4.1.2 Categorical task formulation
Deep neural networks often struggle to match the performance of ML baselines on regression tasks, particularly within imbalanced data regimes [grinsztajn2022treebasedmodelsoutperformdeep, yang2021delvingdeepimbalancedregression]. To address these challenges and to exploit proven model designs for classification tasks, we reformulate precipitation regression as a classification problem. Given a target regression label $y_{\text{reg}}\in\mathbb{R}$ and a training dataset $\mathcal{D}$, we approximate categorical equivalent labels by partitioning the target space into $n$, non-overlapping sub-intervals with step size $\delta$:
@@
We then generate one-hot categorical labels $y\in\{0,1\}^{n}$, representing a bin centered at $\mathcal{D}_{\text{ymin}}+i\delta$, by calculating the nearest label-bin to a target: $i=\text{round}(\frac{y_{\text{reg}}-\mathcal{D}_{\text{ymin}}}{\delta})$. Our model learns to predict categorical labels derived above. Taking the output of the last transformer encoder block, we splice out the ultimate class token $\mathbf{z}^{(L)}_{(cls)}$; a one-layer MLP prediction head generates class-probabilities using this token as input.
$\hat{y}=\text{softmax}\big(\text{MLP}(\mathbf{z}^{(L)}_{(cls)})\big)$ (7)
@@ 4.1.3 Class-weighted cross-entropy loss
$\mathcal{L}(y,\hat{y})=-\sum_{i=1}^{n}w_{i}\log(\hat{y}_{i})y_{i}$ (8)
[…]
@@ 5.5 Ablation study
We conduct ablations for several design choices made throughout this paper. Concretely, our SaTformer can be derived step-by-step from a TimeSformer baseline [bertasius2021spacetimeattentionneedvideo] implemented with interleaved space-time attention modules and trained via standard CCE loss. As previously mentioned, datasets for extreme weather forecasting commonly suffer from large class imbalances. As shown in Figure 2 our target distribution is heavily skewed towards low/no rain events and contains large gaps in sample coverage at the extremes. To encourage balanced performance across classes we implement the weighted CCE loss proposed in equation (8); ablations performed using bin-averaged metrics confirm that including a class-frequency prior from the training distribution greatly aids model skill (Table 1).

### v2

@@ 1 Introduction
Extreme precipitation poses outsized risks to life, property, and commerce. Recent studies predict that these events are likely to increase in frequency and severity during the coming decades, spurred on by a warming atmosphere Madakumbura et al. (2021); Kotz et al. (2022). Timely and accurate public advisories save lives and allow interested parties to make appropriate preparations for future weather events. However, traditional sources of operational forecast guidance are plagued by key limitations. First, numerical weather prediction (NWP) models like the operational ECMWF incur enormous computational costs. And second, these models are often only available at hourly time-scales, and are therefore inappropriate sources of real-time guidance.
Recent data-driven, so-called AI-weather prediction (AI-WP) approaches offer a compute-efficient alternative to NWP models by learning to emulate forecast analysis data. Breakthrough works from Google Lam2023; Price2024, NVIDIA pathak2022fourcastnetglobaldatadrivenhighresolution, Huawei Bi_2023, and others replace core components of the weather-prediction pipeline with neural networks abdi2025hrrrcastdatadrivenemulatorregional; Flora2025; Das2024; Asperti2024; Reulen2024; Ha2024. These learning-based methods have manifold advantages. Namely, AI-WP pipelines can practically run at sub-hourly intervals and in large ensembles; researchers are therefore optimistic that operational meteorologists can one day leverage these models as a source of real-time guidance and robust uncertainty estimation. However, previous studies primarily explore AI-WP for medium-range forecasting on a global scale. This is a natural choice given the surplus of historical data available at these resolutions. By contrast, forecasting extreme precipitation at nowcasting timescales is a nascent area of research.
@@ 2.1 Weather4Cast competition
Beginning in 2021, the annual Weather4Cast competition (W4CC) evaluates machine learning (ML) techniques for weather forecasting within spatial and temporal transfer learning regimes pmlr-v220-gruca23a; gruca2024multitask. Each year, participants are tasked with developing approaches to predict unseen (OPERA) high-resolution radar precipitation fields (or targets derived from these data) using 11-band HRIT geostationary satellite data and static topographic maps as context. Competition data cover 10, non-overlapping regions (spatial areas of $\times1512$ km) across Europe, northern Africa, and the middle-east for which HRIT data are completely available, while OPERA data are sparse.
@@ 3.1 W4CC 2021
In its inaugural year organizers proposed a multivariate weather forecasting challenge: participants predicted future atmospheric states (represented by four variables) for eight hours of lead-time using the same features and topographic information as model inputs kwok2021variationalunetweatherforecasting. The authors of kwok2021variationalunetweatherforecasting proposed to solve this task by training a U-Net VAE with a combination of deterministic and probabilistic losses. In this work it is argued that these design choices are crucial for modeling the fundamental uncertainty (i.e., aleatoric and epistemic) present in weather forecasting and that they therefore aid model performance. However, this complex, multi-loss setup may prove difficult to extend to new tasks in practice.
@@ 3.2 W4CC 2022
In belousov2022solvingweather4castchallengevisual an ensemble of video classification models (e.g., Swin UNETR, ViViT) is presented hatamizadeh2022swinunetrswintransformers; arnab2021vivitvideovisiontransformer). An obvious drawback is that this approach requires training three different models from scratch for a single task and incurs further computational costs by bootstapping predictions at inference-time. Additionally, modifications made by the authors to adopt video classification models to a video generation task (e.g., appending a single convolutional layer to ViViT to predict the output temporal dimension) are likely brittle design choices. In li2022superresolutionprobabilisticrainprediction, the authors apply a 3D U-Net and an Earthformer çiçek20163dunetlearningdense; gao2023earthformerexploringspacetimetransformers to solve this year’s stage 1 and stage 2 challenges, respectively. In the second challenge, an Earthformer is employed to efficiently decompose a video-sequence into cuboids (as opposed to frame-wise patches as proposed in ViT); self-attention is performed between cuboids and a collection of global context tokens $\mathcal{G}$. However, inter-cuboid attention is axial, and therefore, cuboids only learn spatial dependencies from a set of coarse global context vectors.
Beginning with a standard 3D U-Net, the authors of park2022rainunetsuperresolutionrainmovie introduce a time-space (TS) convolutional block in which a spatial convolution, followed by a “temporal-wise dilated convolution” (TSDConv), and finally a temporal convolution are applied to the input feature maps. The authors note that their proposed TSDConv is simply an extension of the large kernel convolution operation into three dimensions. Additionally, WeatherFusionNet pihrt2022weatherfusionnetpredictingprecipitationsatellite propose a physically informed U-Net-based architecture. Their model as inputs uses HRIT satellite radiances, a PDE-constrained video prediction via PhyDNet guen2020disentanglingphysicaldynamicsunknown, and the outputs of a proposed Sat2Rad module to forecast binary rain masks. While the methods above achieve strong performance on W4CC tasks, most are fundamentally reliant on CNN model designs, and likely suffer from the inductive biases and poor scaling these architectures exhibit in other domains.
@@ 3.3 W4CC 2023
This year, researchers at Alibaba Cloud proposed a mixture-of-experts-inspired-solution to both the “core” and “transfer learning” competition tracks li2023precipitationpredictionusingensemble. The authors design a “predictor”, which samples multiple times from a WeatherFusionNet backbone, and a “controller”; the latter module produces a final output as a weighted combination of the expert predictions. In addition to being a minor extension of prior works, the proposed model employs an arduous three-stage training scheme. The aforementioned work is still preferable to kumar2023skilfulprecipitationnowcastingusin, who simply finetune a NowcastNet Zhang2023 on the challenge dataset. In han2023learningrobustprecipitationforecaster an augmentation technique is proposed to sample linear interpolations of the original training samples. By formulating precipitation forecasting as a multi-class, video classification task (where each per-pixel class represents a discrete interval of precipitation, the authors may employ a modified dice-loss with an added ordinal prior (i.e., under-predictions are punished more heavily than over-predictions). While certainly performant, this approach is heavily engineered and therefore may generalize poorly to new tasks and datasets
@@ 3.4 W4CC 2024
deshpande2024conditionalgenerativeadversarialnetwork develop a conditional GAN network trained on a sparsely-sampled subset of the original HRIT channels. Here, the authors completely discard near-infrared, water vapor, and visible satallite data, and average the remaining 4 channels; after data preprocessing their model is trained on only single-channel inputs. Further preprocessing techniques are employed to improve model performance: including thresholding out non-cloudy regions and extracting optical flow scores from the input sequence. Similar to han2023learningrobustprecipitationforecaster, this work proposes a bespoke, highly-engineered pipeline that is unlikely to extend gracefully to future work.
@@ 4.1 SaTformer model
Following prior work dosovitskiy2021imageworth16x16words; bertasius2021spacetimeattentionneedvideo, we partition each frame in the input sequence into $N=HW/P^{2}$ non-overlapping patches with side-length $P$ indexed at timestep $t\in 1,...,T$ and spatial index $p\in 1,...,N$: $\mathbf{x}^{(0)}_{(t,p)}\in\mathbb{R}^{P\times P\times C}$; patches completely cover the spatial dimensions of the original frame. We flatten each patch, then project the resulting vectors into tokens: $\mathbf{z^{\prime}}^{(0)}_{(t,p)}\in\mathbb{R}^{d}$. We prepend a randomly initialized “class token” $\mathbf{z^{\prime}}_{(0,0)}^{(0)}$ to the token sequence intended to aggregate relevant information for prediction of the target label:
$\mathbf{z^{\prime}}_{(0,0)}^{(0)},\mathbf{z^{\prime}}^{(0)}_{(0,1)},\dots,\mathbf{z^{\prime}}^{(0)}_{(0,N)},\mathbf{z^{\prime}}^{(0)}_{(1,0)},\dots,\mathbf{z^{\prime}}^{(0)}_{(T,N)}]$ (1)
@@ 4.1.1 Full space-time self-attention
Where $\mathbf{q}^{(l)}_{(t,p)}\in\mathbb{R}^{d}$, $\mathbf{k}^{(l)}_{(t,p)}\in\mathbb{R}^{d}$ $\mathbf{v}^{(l)}_{(t,p)}\in\mathbb{R}^{d}$ and $\text{LN}(\cdot)$ denotes the LayerNorm operation ba2016layernormalization. Next we compute importance weights $\mathbf{a}^{(l)}_{(t,p)}$ between all pairs of tokens across time and space:
$\mathbf{a}^{(l)}_{(t,p)}=\text{softmax}\big(\frac{(\mathbf{q}^{(l)}_{(t,p)})^{\top}}{\sqrt{d}}\cdot[\mathbf{k}_{(0,0)}^{(l)},\mathbf{k}^{(l)}_{(0,1)},\dots,\mathbf{k}^{(l)}_{(0,N)},\mathbf{k}^{(l)}_{(1,0)},\dots,\mathbf{k}^{(l)}_{(T,N)}]\big)$ (4)
@@
$\mathbf{s}^{(l)}_{(t,p)}=\mathbf{a}^{(l)}_{(t,p)(0,0)}\mathbf{v}^{(l)}_{(0,0)}+\sum_{t^{\prime}=1}^{T}\sum_{p^{\prime}=1}^{N}\underbrace{\mathbf{a}^{(l)}_{(t,p)(t^{\prime},p^{\prime})}}_{\text{scalar}}\mathbf{v}^{(l)}_{(t^{\prime},p^{\prime})}$ (5)
@@
$\mathbf{z^{\prime}}^{(l)}_{(t,p)}=\mathbf{s}^{(l)}_{(t,p)}W_{\text{out}}+\mathbf{z}^{(l-1)}_{(t,p)}$ $\mathbf{z}^{(l)}_{(t,p)}=\text{MLP}(\text{LN}(\mathbf{z^{\prime}}^{(l)}_{(t,p)}))+\mathbf{z^{\prime}}^{(l)}_{(t,p)}$ (6)
@@ 4.1.2 Categorical task formulation
Deep neural networks often struggle to match the performance of ML baselines on regression tasks, particularly within imbalanced data regimes grinsztajn2022treebasedmodelsoutperformdeep; yang2021delvingdeepimbalancedregression. To address these challenges and to exploit proven model designs for classification tasks, we reformulate precipitation regression as a classification problem. Given a target regression label $y_{\text{reg}}\in\mathbb{R}$ and a training dataset $\mathcal{D}$, we approximate categorical equivalent labels by partitioning the target space into $n$, non-overlapping bins spaced by step size $\delta$:
@@
We then generate one-hot categorical labels $y\in\{0,1\}^{n}$, representing a bin centered at $\mathcal{D}_{\text{ymin}}+i\delta$, by calculating the nearest label-bin to a target: $i=\text{round}(\frac{y_{\text{reg}}-\mathcal{D}_{\text{ymin}}}{\delta})$. Our model learns to predict categorical labels derived above. Taking the output of the last transformer encoder block, we splice out the ultimate class token $\mathbf{z}^{(L)}_{(0,0)}$; a one-layer MLP prediction head generates class-probabilities using this token as input.
$\hat{y}=\text{softmax}\big(\text{MLP}(\mathbf{z}^{(L)}_{(0,0)})\big)$ (7)
@@ 4.1.3 Class-weighted cross-entropy loss
$\mathcal{L}(y,\hat{y})=-\sum_{i=0}^{n-1}w_{i}\log(\hat{y}_{i})y_{i}$ (8)
Here $\hat{y}$ is our model’s predicted probability distribution over the target classes and $w_{i}=-\log(\frac{|\mathcal{D}_{i}|}{|\mathcal{D}_{\text{total}}|})$ is the log-scaled relative frequency of a class in the training set.
@@ 5.5 Ablation study
We conduct ablations for several design choices made throughout this paper. Concretely, our SaTformer can be derived step-by-step from a TimeSformer baseline bertasius2021spacetimeattentionneedvideo implemented with interleaved space-time attention modules and trained via standard CCE loss. As previously mentioned, datasets for extreme weather forecasting commonly suffer from large class imbalances. As shown in Figure 2 our target distribution is heavily skewed towards low/no rain events and contains large gaps in sample coverage at the extremes. To encourage balanced performance across classes we implement the weighted CCE loss proposed in equation (8); ablations performed using bin-averaged metrics confirm that including a class-frequency prior from the training distribution greatly aids model skill (Table 1).